y= -3/4x - 10
how do I find the slope and y intercept?

Answer by stanbon(75887)

y= -3/4x - 10
how do I find the slope and y intercept?
---
The slope-intercept form is y = mx + b
where m is the slope and b is the y-intercept.
----
Since your equation is already in that form,
m = -3/4 is the slope
---
and b = -10 is the y-intercept.
===================================
Cheers,
Stan H.
==========